# L4P: Low-Level 4D Vision Perception Unified

### v1

@@ 3.3 Sparse Prediction Heads
Given a pixel prompt, $(t_{i},x_{i},y_{i})$, in a video, we want to estimate the corresponding 3D trajectory, $\mathcal{T}_{i}=\{\hat{x}_{i}(t),\hat{y}_{i}(t),\hat{d}_{i}(t),\hat{v}_{i}(t)\}_{t=0}^{S-1}$, where at time $t$, $(\hat{x}_{i}(t),\hat{y}_{i}(t))$ denotes the 2D track location, $\hat{d}_{i}(t)$ is the track depth with respect to camera, and $\hat{v}_{i}(t)$ is the track visibility indicating if a track is visible or occluded. This is a challenging task since it requires tracking the pixel in 2D when visible, tracking it through occlusions, and reasoning about the depth of the track. Moreover, our video encoder has limited temporal context, since it can only process videos with fixed temporal window of $T$ frames, and we want to enable tracking for arbitrarily long videos with $S>T$ frames. This makes adapting a general-purpose head particularly challenging. To tackle this, we adapt an online approach. We propose a head that allows us to estimate the 3D track for an input pixel prompt within the temporal context ($T$ frames) of the video encoder. For online estimation beyond $T$ frames, we propose a memory mechanism for our head and a recipe to train it efficiently (see Figure 3 for an overview).
@@ 3.3.1 Tracking within the Temporal Context
Posing the sparse tracking task within our unified framework requires special care. Instead of directly estimating point-track positions, we propose to represent tracks using dense probability heatmaps. Casting tracking as a problem of estimating pixel-aligned 2D maps affords us a shared representation between sparse and dense tasks, which is critical for using a shared backbone. To achieve this, we adapt the prompt-encoding and mask-decoding mechanisms from SAM [28]. The input pixel prompt is encoded using 3D positional encoding and a learnable embedding to generate input point token $\mathcal{P}$ with embedding dimension of $C$. Similarly, we define output tokens with learnable embeddings to estimate different components of a 3D track: a heatmap token ($\mathcal{H}$) to estimate the 2D pixel position of the track across the video, a depth ($\mathcal{D}$) and a visibility ($\mathcal{V}$) token. Input and output tokens interact with the video tokens, $\mathcal{S}$, also encoded using 3D positional encoding, using a two-way attention mechanism to decode the video features. These video features are then reshaped and up-sampled, and a final inner product with the processed output tokens gives us output masks of size $T\times H\times W$. For the 2D track estimation, we interpret this output mask as a probability heatmap that encodes the 2D track position and apply a 2D soft-argmax operation to estimate the 2D track position $(\hat{x}_{i}(t),\hat{y}_{i}(t))$ at each frame $t$. For depth and visibility, we simply apply a 2D average pooling operation, followed by exponential and sigmoid operations respectively to estimate the track depth $\hat{d}_{i}(t)$ and the visibility $\hat{v}_{i}(t)$ at each frame $t$. This simple design also allows us to query points anywhere in the video and track them in parallel. We adapt two-way attention from SAM [28] and keep this head lightweight by using only two two-way attention layers. We also replace the 2D convolutions in the original mask-decoder of SAM with 3D convolutions to enable temporal reasoning.
@@ 3.3.2 A Memory Mechanism for Long Videos
To track beyond a single window of length $T$ frames, we adopt an online approach.
@@
A naïve approach is to chain tracks across windows. Given two consecutive and overlapping windows, and a 2D track estimated in the first one, we can use a point on the track in the temporal overlap between the two windows as the query for tracking in the second one. To pick a good point to chain the tracks, we can select the one with highest visibility score. However, this solution is brittle. First, a tracked point may not be visible in the overlap between the windows. To tackle this problem, inspired by Karaev et al. [26], we introduce a track-feature token $\mathcal{F}$ that is passed to subsequent windows as an additional prompt (see Figure 3). However, unlike Karaev et al., we do not initialize it explicitly with the local appearance around the query point, so the two-way attention head is free to capture the most useful information to track through occlusions. Second, the naïve solution described above does not allow the system to reason across temporal windows, which makes it prone to drifting or to losing tracks. The track-feature tokens help, but to provide even more cross-window information, we pass the video tokens decoded by the two-way attention stage of the current window to the next, as shown in Figure 3. We achieve this by projecting the decoded video tokens in the overlapping region via a linear layer, and by adding them to the corresponding incoming video tokens to the two-way attention stage in the next window. Our memory strategy based on these two mechanisms is critical to allow proper reasoning across temporal windows as shown by the comparison in ablation study in Section 4.5.
@@
Online training. Training the memory mechanism requires unrolled-window training [26], in which we compute the video features for all the overlapping windows in a video of length $S$, and then compute the tracks for the entire video in an online fashion. However, training such an approach end-to-end is prohibitive due to memory constraints. To alleviate this, we adapt a two-stage training strategy. First, we train only for a single window but train all the parameters of our network, and in a second stage we freeze all but the last few layers of our video encoder, and fine-tune it along with the tracking head for unrolled window training. In Tables 3 and 4 we show this approach improves the performance over the naïve chaining approach.
@@ 4 Experiments
First, we provide details about our architecture and training. Then, for each task, we discuss the baselines, the evaluation metrics and datasets, and show quantitative and qualitative results. We also present an ablation study that evaluates the contribution of different components of our approach. Finally, we show a way to extend our approach to new tasks by using motion-based segmentation as an example.
4.1 Implementation
Training datasets. We use the video encoder from Wang et al. [58], which is pre-trained on 1.35M video clips for masked auto-encoding. To fine-tune our model, we use a limited number of synthetic datasets covering a varying range of types of annotations, and rely on the priors from the pre-trained video encoder for generalization. We use Kubric [16], a synthetic dataset in which multiple objects interact, annotated with depth, flow, and 2D and 3D point tracking. To include videos with long 3D trajectories, we add PointOdyssey [77] and DynamicReplica [25]. Both are synthetic datasets with animated characters in mostly indoor scenes. Both datasets have depth annotations and, in addition, DynamicReplica offers optical flow annotations. To further increase scene diversity, we also include TartanAir [62], which provides annotations for flow and depth.
Architectures. Our video encoder [58] processes video-clips of size $16\times 224\times 224$. It uses a patch size of $2\times 14\times 14$, which results in $P=2048$ video tokens, and an embedding dimension of $C=1408$. It has 40 encoder blocks, and we use the output from blocks 14, 21, 28, 36 for DPT heads for dense tasks, while the sparse heads use features only from the last block. Feeding the sparse and dense heads with tokens from different blocks allows us to maintain the performance on dense tasks while we fine-tune our model to train the memory mechanism for the tracking tasks, as we discuss below. For a $16\times 224\times 224$ video clip, our method generates the outputs for all our tasks in $\sim$300ms on an NVIDIA A6000 GPU. This corresponds to $\sim$19ms for a single frame, which is competitive with single-task approaches (see Supplementary for detailed comparisons). However, our method’s latency may prevent its use for applications that require strict real-time performance.
Training. We initialize our video encoder using a pretrained VideoMAE [58] and fine-tune our model in two stages. In both stages, we construct a batch of many tracks per video for the tracking task. In the first stage, we train end-to-end for depth, flow, 2D and 3D point tracking tasks on a single window of $T=16$ frames. In the second stage, we further fine-tune our model for the tracking tasks using unrolled-window training and the memory mechanism for online tracking. We train on videos of length $S=40$ by using 4 overlapping windows of size 16 frames and a stride of 8. Due to memory constraints, in the second stage, we freeze all the parameters, except the last three layers (37-39) of the video encoder and the sparse task head. This allows us to maintain the performance on depth and flow, while training the memory mechanism to improve the tracking tasks. Both stages use a batch size of 8 and are trained on a single 8-GPU (NVIDIA A100) node for 100k iterations. Training takes 1 day and 2 days respectively.
Losses. We use the SILog [11] loss for depth and L1 loss for optical flow. For tracking, we use L1 loss for 2D track positions, scale-invariant loss for track depth (similar to dense depth), and binary cross entropy loss for track visibility. Like with the choices of tasks heads, we pick the most widely used losses for each of our tasks. However, since we train for multiple tasks at once, weighting the losses appropriately is critical. We find the loss weights empirically by first bringing the losses in the same order of magnitude and then doing a small hyperparameter search around those weights.
Please refer to the Supplementary for additional implementation details.
@@ 4.1 Implementation
DepthCrafter [20] $0.270$ $\mathbf{0.697}$ $0.123$ $0.856$ $0.104$ $0.896$ $0.071$ $\underline{0.972}$ $0.072$ $0.948$
L4P-depth (Ours) $\underline{0.251}$ $0.659$ $0.102$ $0.895$ $0.099$ $0.916$ $0.061$ $\underline{0.972}$ $0.078$ $0.932$
L4P-depth* (Ours) $0.267$ $\underline{0.693}$ $\mathbf{0.070}$ $\mathbf{0.954}$ $0.097$ $0.903$ $\mathbf{0.057}$ $\underline{0.972}$ $0.078$ $0.932$
L4P (Ours) $0.263$ $0.662$ $0.103$ $0.898$ $\underline{0.093}$ $\underline{0.924}$ $0.060$ $\mathbf{0.973}$ $0.081$ $0.925$
L4P * (Ours) $\mathbf{0.247}$ $0.691$ $\underline{0.072}$ $\underline{0.951}$ $\mathbf{0.090}$ $\mathbf{0.928}$ $\underline{0.058}$ $\mathbf{0.973}$ $0.081$ $0.925$
Table 1: Zero-shot depth estimation results. We compare our methods against both single-image baselines (Row 1-3) and SOTA video depth estimation approaches (Row 4-6). L4P-depth refers to our model trained specifically for depth estimation and L4P refers to the version trained jointly for all tasks. Models marked * have predictions in overlapping windows aligned using the strategy described in Section 3.2. On video datasets (all except NYUv2), our model consistently performs better than DepthCrafter, the closest competition, and by a large margin on ScanNet and KITTI. Best and second best results are highlighted.
Figure 5: Qualitative results for depth estimation. We include one example each from Bonn, KITTI, and ScanNet. Inference is conducted on 16-frame clips, but only 1 frame is shown.
@@ 4.2 Video Depth Estimation
We follow DepthCrafter [20] and evaluate video depth estimation on a collection of five datasets. We do not use any of the datasets for training our models or the baselines to better understand their generalization abilities. There is an inherent scale-ambiguity in the estimated depthmaps. We follow the common practice of aligning linearly the estimation with the GT before calculating evaluation metrics. The alignment is done for all the frames at once, and is carried out in disparity space via least-square fitting. For comparison on single image datasets, we repeat the single frame 16 times to compute our estimations. We report two metrics: AbsRel ($\text{mean}(|\hat{\mathbf{d}}-\mathbf{d}|/\mathbf{d}))$ and $\delta_{1}$ (ratio of pixels satisfying $\max(\mathbf{d}/\hat{\mathbf{d}},\hat{\mathbf{d}}/\mathbf{d})<1.25$), where $\mathbf{d}$ represents GT, and $\hat{\mathbf{d}}$ is depth estimation after alignment. We upsample our estimations from $224\times 224$ to each dataset’s original resolution for evaluation.
We consider video approaches including NVDS [63], ChronoDepth [48], DepthCrafter [20], as well as single-image ones, including Marigold [27] and DepthAnything [73, 74]. Among them, DepthCrafter [20] and DepthAnything [73, 74] each represent the SOTA respectively. Marigold and DepthCrafter are diffusion models, which afford impressive levels of details, but require an expensive iterative denoising process.
Our results show consistent advantages over both SOTA single-image and video depth approaches on the four video datasets (Table 1). Since L4P is a video approach, applying it on single images from NYUv2 does not provide the necessary temporal context for it to perform well. DepthCrafter also similarly suffers on NYUv2. Figure 5 shows qualitative samples and comparison with select SOTA approaches. L4P produces a level of details comparable to that of diffusion models such as DepthCrafter, while generally capturing more accurate relative scales.
Discussion. Our final model performs on par with a specialized depth model (Table 1), despite optimized jointly for all of our tasks. Scale alignment between windows for online inference makes a significant impact on ScanNet. This is due to the fast-paced view change in ScanNet samples making scale inconsistency between windows more prominent. It is also worth noting that L4P performs competitively on KITTI, despite not fine-tuned on synthetic datasets that include driving scenarios.
@@
Ours $(0.13)$ $(97.6)$ $(0.03)$ $(99.9)$ $\mathbf{0.10}$ $\mathbf{98.5}$
Table 2: Optical flow estimation results. We evaluate on validation sets of all datasets. Our model has seen Kubric and Dynamic Replica during training (numbers in brackets). All others show cross-dataset generalization. On Spring, we surpass both two-frame (marked with *) and multi-frame SOTA baselines, despite the latter having specially designed architectures and a complex memory mechanism.
Figure 6: Qualitative results for optical flow estimation on Spring. Our results compare favorably to baselines in terms of both details and accuracy. Inference is conducted on 16-frame clips, but only 1 frame is shown.
@@ 4.3 Multi-Frame Optical Flow Estimation
We consider two baselines for comparison. RAFT [53], a competitive and widely used two-frame approach, creates dense pairwise pixel features and uses recurrent updates to estimate optical flow. MemFlow [10], a recently published work, ranks among the top methods on the Spring benchmark. It is a multi-frame approach that relies on a memory mechanism to leverage temporal context. Quantitatively, L4P compares favorably to both RAFT and MemFlow on Spring (Table 2). Our model can capture well both small and large motions and presents more precise motion boundaries (Figure 6). In addition, multi-frame approaches like MemFlow and ours generally have an edge in temporal stability (see Supplementary). Unlike many specialized approaches, our model currently only operates on low-resolution videos and further work is needed to enable efficient high-res estimation.
@@
Ours (2D Only) $\mathbf{56.7}$ $54.2$ $49.8$ $53.5$ $69.4$ $88.6$
Ours (w/o Mem) $36.8$ $47.4$ $41.1$ $41.8$ $62.9$ $78.6$
Ours $53.0$ $51.6$ $48.8$ $51.2$ $67.0$ $\mathbf{88.7}$
Table 3: Evalution of 2D point tracking on TAPVid-3D. 2D GT trajectories are obtained by projecting 3D GT trajectories onto 2D. Though behind 2D SOTA approaches, our model performs competitively once trained specifically for 2D tracking (“2D Only”).
@@
Ours (w/o Mem) 8.2 15.4 72.7 5.5 10.0 83.3 $15.1$ $25.2$ $79.9$ $9.6$ $16.9$ $78.6$
Ours $\mathbf{11.2}$ $\mathbf{17.7}$ $\mathbf{90.3}$ $6.6$ $11.4$ $\mathbf{88.1}$ $\mathbf{18.6}$ $\mathbf{28.2}$ $\mathbf{87.6}$ $\mathbf{12.1}$ $\mathbf{19.1}$ $\mathbf{88.7}$
Table 4: Evaluation of 3D tracking on the full_eval split of TAPVid-3D. The top approaches combine 2D point tracking approaches with COLMAP (CM) [45], while the bottom ones, including ours are feedforward. Our approach consistently outperforms previous feedforward works, and also COLMAP baselines on average. We also show the impact of our memory mechanism (Ours vs. Ours w/o Mem). “ZD” refers to ZoeDepth.
Figure 7: Qualitative results of Sparse 2D/3D tracking on the TAPVid-3D benchmark. Comparison with SpaTracker, a SOTA 3D tracking approach, demonstrates the superior quality of our 2D and 3D tracks. For joint visualization of depth and 3D tracks, we align them using median scaling. We use our depth maps for visualization of GT and for SpaTracker we use the ones used by their approach.

### v2

@@ 3.3 Sparse Prediction Heads
Given a video and pixel prompt, $(t_{i},x_{i},y_{i})$, we want to estimate the corresponding 3D trajectory, $\mathcal{T}_{i}=\{\hat{x}_{i}(t),\hat{y}_{i}(t),\hat{d}_{i}(t),\hat{v}_{i}(t)\}_{t=0}^{S-1}$, where at time $t$, $(\hat{x}_{i}(t),\hat{y}_{i}(t))$ denotes the 2D track location, $\hat{d}_{i}(t)$ is the track depth with respect to camera, and $\hat{v}_{i}(t)$ is the track visibility indicating if a track is visible or occluded at time $t$. This is a challenging task since it requires tracking the pixel in 2D when visible, tracking it through occlusions, and reasoning about its depth along the track. Moreover, our video encoder has limited temporal context, since it can only process videos with fixed temporal window of $T$ frames, and we want to enable tracking for arbitrarily long videos with $S>T$ frames. This makes adapting a general-purpose head particularly challenging. To tackle this, we introduce an online approach. We design a head that allows us to estimate the 3D track for an input pixel prompt within the temporal context ($T$ frames) of the video encoder. For online estimation beyond $T$ frames, we propose a memory mechanism for our head and a recipe to train it efficiently (see Figure 3).
@@ 3.3.1 Tracking within the Temporal Context
Posing the sparse tracking task within our unified framework requires special care. Instead of directly estimating point-track positions, we propose to represent tracks using dense probability heatmaps. Casting tracking as a problem of estimating pixel-aligned 2D maps affords us a consistent representation between sparse and dense tasks, which is critical for using a shared backbone. To achieve this, we build on the prompt-encoding and mask-decoding mechanisms from SAM [29]. We encode the input pixel prompt using 3D positional encoding and a learnable embedding to generate input point token $\mathcal{P}$ with embedding dimension of $C$. Similarly, we define output tokens with learnable embeddings to estimate different components of a 3D track: a heatmap token ($\mathcal{H}$) to estimate the 2D pixel position of the track across the video, a depth token ($\mathcal{D}$), and a visibility token ($\mathcal{V}$). Input and output tokens interact with the video tokens, $\mathcal{S}$, also encoded with 3D positional encoding, using a two-way attention mechanism to decode the video features. These video features are then reshaped and upsampled, and a final inner product with the processed output tokens gives us output masks of size $T\times H\times W$. For 2D track estimation, we interpret this output mask as a probability heatmap that encodes the 2D track position and apply a 2D soft-argmax operation to estimate the 2D track position $(\hat{x}_{i}(t),\hat{y}_{i}(t))$ at each frame $t$. For depth and visibility, we simply apply a 2D average pooling operation, followed by exponential and sigmoid operations respectively to estimate the track depth $\hat{d}_{i}(t)$ and the visibility $\hat{v}_{i}(t)$ at each frame $t$. This design also allows us to query points anywhere in the video and track them in parallel. We use two lightweight, two-way attention layers [29]. We use 3D convolutions to enable temporal reasoning.
@@ 3.3.2 A Memory Mechanism for Long Videos
To track beyond a single window of length $T$ frames, we propose an online strategy. A naïve approach would be to chain tracks across windows. Given two consecutive and overlapping windows, and a 2D track estimated in the first one, we can use a point on the track in the temporal overlap between the two windows as the query for tracking in the second one. To pick a good point to chain the tracks, we can select the one with highest visibility score. However, this solution is brittle for two reasons. First, a tracked point may not be visible in the overlap between the windows. To tackle this problem, inspired by Karaev et al. [27], we introduce a track-feature token $\mathcal{F}$ that is passed to subsequent windows as an additional prompt (see Figure 3). However, unlike Karaev et al., we do not initialize it explicitly with the local appearance around the query point, so the two-way attention head is free to capture the most useful information to track through occlusions. Second, the naïve solution described above does not allow the system to reason across temporal windows, which makes it prone to drifting or to losing tracks. The track-feature tokens help, but to provide even more cross-window information, we pass the video tokens decoded by the two-way attention stage of the current window to the next, as shown in Figure 3. We achieve this by projecting the decoded video tokens in the overlapping region via a linear layer, and by adding them to the corresponding incoming video tokens to the two-way attention stage in the next window. Our memory strategy based on these two mechanisms is critical to allow proper reasoning across temporal windows as shown by the comparison in ablation study in Section 5.4.
@@
Online training. Training the memory mechanism requires unrolled-window training [27], in which we compute the video features for all the overlapping windows in a video of length $S$, and then compute the tracks for the entire video in an online fashion. However, training such an approach end-to-end is prohibitive due to memory constraints. To alleviate this, we adapt a multi-stage training strategy. First, we train only for a single window but train all the parameters of our network. In the next stage, we freeze all but the last few layers of our video encoder, and fine-tune it along with the tracking head for unrolled window training. In Table 5, we show this approach improves the performance over the naïve chaining approach.
4 Implementation
The training curriculum is critical to leverage the VideoMAE priors and fine-tune our system for all tasks at once.
Training datasets. We use the video encoder from Wang et al. [62], which is pre-trained on 1.35M video clips for masked auto-encoding. To fine-tune our model, we use four synthetic datasets covering various types of annotations, and rely on the priors from the pre-trained video encoder for generalization. We use Kubric [16], a synthetic dataset in which multiple objects interact, annotated with ground truth (GT) depth, flow, and 2D/3D point tracking. To include videos with long 3D trajectories, we add PointOdyssey [83] and DynamicReplica [26]. Both are synthetic datasets with animated characters in mostly indoor scenes. Both have depth GT and DynamicReplica also offers optical flow GT. To further increase scene diversity, we also include TartanAir [68], which provides GT for flow and depth.
Architectures. Our video encoder processes $16\times 224\times 224$ clips. It uses a patch size of $2\times 14\times 14$, which results in $P=2048$ video tokens, and an embedding dimension of $C=1408$. It has 40 encoder blocks, and we use the output from blocks 14, 21, 28, 36 for DPT heads for dense tasks, while the sparse heads use features only from the last block (block 39). Feeding the sparse and dense heads with tokens from different blocks is critical for maintaining the performance on dense tasks while we fine-tune our model to train the memory mechanism for the tracking tasks, as we discuss below. For a $16\times 224\times 224$ video clip, our method generates the outputs for all our tasks, including the additional tasks, in $\sim$300ms on an NVIDIA A6000 GPU. This corresponds to $\sim$19ms for a single frame, which is competitive with single-task approaches (see Supplementary for detailed comparisons). However, our method’s latency may prevent its use for applications that require strict real-time performance.
Training. Perhaps unsurprisingly, training our system to maximize the performance on all tasks at once requires particular care. We initialize the video encoder using a pre-trained VideoMAE [62] and fine-tune the complete system in three stages. All the stages use a batch size of 8 and are trained on a single 8-GPU (NVIDIA A100) node for 100k iterations. In all the stages, we construct a batch of multiple tracks per video for the tracking task. In the first stage, we train end-to-end for depth, flow and point tracking tasks on a single window of $T=16$ frames and train only on Kubric. In the second stage of training, we add the remaining three datasets and further fine-tune the model on a single window for all the tasks. Since not all the datasets offer GT for all tasks, we form our batches by sampling two videos from each of the four datasets. This forces each batch to provide training signal for all task, and thus prevents training instability. In the third stage, we further fine-tune our model for the tracking tasks using unrolled-window training and our memory mechanism for online tracking. We train on videos of length $S=40$ by using 4 overlapping windows of size 16 frames and a stride of 8. Due to memory constraints, in the third stage, we freeze all the parameters, except the last three layers (37-39) of the video encoder and the sparse task head. This allows us to maintain the performance on depth and flow, while training the memory mechanism to improve the tracking tasks. Training takes around 4 days for all the stages combined.
Losses. We use the SILog [11] loss for depth and L1 loss for optical flow. For tracking, we use L1 loss for 2D track positions, scale-invariant loss for track depth (similar to dense depth), and binary cross entropy loss for track visibility. Like with the choices of tasks heads, we pick the most widely used losses for each of our tasks. However, since we train for multiple tasks at once, weighting the losses appropriately is critical. We find the loss weights empirically by bringing the losses in the same order of magnitude and then doing a small hyperparameter search around those weights.
Please refer to the Supplementary for additional important implementation details.
@@ 5 Experiments
Here we discuss our results and ablate the contribution of different components of our approach. We also show that L4P can be extended to new tasks with no performance degradation for the existing ones.
@@
DepthCrafter [21] $0.270$ $0.697$ $0.123$ $0.856$ $0.104$ $0.896$ $0.071$ $0.972$ $0.072$ $0.948$
L4P $\mathbf{0.219}$ $\mathbf{0.700}$ $\mathbf{0.071}$ $\mathbf{0.953}$ $\mathbf{0.084}$ $\mathbf{0.935}$ $\mathbf{0.056}$ $\mathbf{0.973}$ $0.084$ $0.916$
Table 1: Depth estimation results. We compare our full model with single-image (Row 1-3) and video depth estimation baselines (Row 4-6). On video datasets (all except NYUv2), our model consistently performs better than DepthCrafter, the closest competition, and by a large margin on ScanNet and KITTI.
Figure 5: Comparisons with SOTA methods for different tasks. We run inference on video clips and show only 1 frame for the top three tasks. Purple overlays and dotted trajectories represent GT for motion-based segmentation and camera pose estimation.
@@ 5.1 Video Depth Estimation
We follow DepthCrafter [21] and evaluate video depth estimation on a collection of five datasets. We do not use any of the datasets for training our models or the baselines to better understand their generalization abilities. When estimating depth for videos longer than $T=16$ frames, we follow the online inference and alignment strategies explained in Section 3.2. There is an inherent scale-ambiguity in the estimated depthmaps. We follow the common practice of aligning linearly the estimation with the GT before calculating evaluation metrics. The alignment is done for all the frames at once, and is carried out in disparity space via least-square fitting. For comparison on single image datasets, we repeat the single frame 16 times to compute our estimations. We report two metrics: AbsRel ($\text{mean}(|\hat{\mathbf{d}}-\mathbf{d}|/\mathbf{d}))$ and $\delta_{1}$ (ratio of pixels satisfying $\max(\mathbf{d}/\hat{\mathbf{d}},\hat{\mathbf{d}}/\mathbf{d})<1.25$), where $\mathbf{d}$ represents GT, and $\hat{\mathbf{d}}$ is depth estimation after alignment. We upsample our estimations from $224\times 224$ to each dataset’s original resolution for evaluation.
We consider video approaches including NVDS [69], ChronoDepth [49], DepthCrafter [21], as well as single-image ones, including Marigold [28] and DepthAnything [78, 79]. Among them, DepthCrafter [21] and DepthAnything [78, 79] each represent the SOTA respectively. Marigold and DepthCrafter afford impressive levels of details, but their denoising process is expensive.
Our results show consistent advantages over both SOTA single-image and video depth approaches on the four video datasets (Table 1). Since L4P is a video approach, applying it on single images from NYUv2 does not provide the necessary temporal context for it to perform well. DepthCrafter also similarly suffers on NYUv2. Figure 5 (first row) shows qualitative samples and comparison with select SOTA approaches. L4P produces a level of details comparable to that of diffusion models such as DepthCrafter, while generally capturing more accurate relative scales. It is also worth noting that our approach performs competitively on KITTI, despite the fine-tuning not including driving scenarios.
@@
L4P $(0.13)$ $(97.5)$ $(0.03)$ $(99.9)$ $\mathbf{0.09}$ $\mathbf{98.7}$
Table 2: Optical flow estimation results. ($\cdot$) indicates that our method was trained on the training set of that dataset. On Spring, we surpass both two-frame (marked with *) and multi-frame SOTA baselines, despite the latter having a specialized architecture and an advanced memory mechanism.
@@ 5.2 Multi-Frame Optical Flow Estimation
We consider two baselines for comparison. RAFT [55], a competitive and widely used two-frame approach, creates dense pairwise pixel features and uses recurrent updates to estimate optical flow. MemFlow [10], a recently published work, ranks among the top methods on the Spring benchmark. It is a multi-frame approach that relies on a memory mechanism to leverage temporal context. Quantitatively, L4P compares favorably to both RAFT and MemFlow on Spring (Table 2). Our model captures both small and large motions and presents more precise motion boundaries (Figure 5, second row). In addition, multi-frame approaches like MemFlow and ours generally have an edge in temporal stability (see Supplementary). Unlike many specialized approaches, our model currently only operates on low-resolution videos and further work is needed to enable efficient high-res estimation.
@@
L4P (2D Only) $\mathbf{55.5}$ $53.0$ $48.7$ $52.4$ $68.2$ $\mathbf{88.7}$
L4P $52.4$ $50.0$ $48.1$ $50.2$ $66.1$ $\mathbf{88.5}$
Table 3: Evaluation of 2D tracking on TAPVid-3D (full_eval split). 2D GT trajectories are obtained by projecting 3D GT trajectories onto 2D. Though behind 2D SOTA approaches, L4P is competitive once fine-tuned only for 2D tracking (“2D Only”).
@@
L4P $\mathbf{11.1}$ $\mathbf{17.7}$ $\mathbf{90.3}$ $6.0$ $10.5$ $\mathbf{87.8}$ $\mathbf{18.3}$ $\mathbf{28.0}$ $\mathbf{87.5}$ $\mathbf{11.8}$ $\mathbf{18.7}$ $\mathbf{88.5}$
Table 4: Evaluation of 3D tracking on TAPVid-3D (full_eval split). The top approaches combine 2D point tracking approaches with COLMAP (CM) [46], while the bottom ones, including ours are feedforward. “ZD” refers to ZoeDepth. Our approach outperforms previous approaches on average across all the metrics.
Figure 6: Sparse 2D/3D tracking on the TAPVid-3D benchmark. Comparison with SpaTracker, a SOTA 3D tracking approach. For joint visualization of depth and 3D tracks, we align them using median scaling. We use our depth maps for visualization of GT and for SpaTracker we use the ones used by their approach.